# Interpretable Visualizations of Data Spaces for Classification Problems

### v1

@@ A.1 Linear and Kernel Principal Components Analysis
$\ell_{proj}=||\mathbf{X-TP_{TX}}||^{2}/n_{\mathbf{{samples}}}$ (11)
@@
A.2 Linear and Quadratic Discriminant Analysis
@@
$f_{k}(x)=\frac{1}{(2\pi)^{n_{\mathbf{{features}}}/2}|\Sigma_{k}|^{1/2}}e^{-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})}$ (12)
@@
For dimensionality reduction, LDA will identify a linear projection that maximizes the separation between classes. One can consider the $n_{\mathbf{classes}}$ class means as vectors in $\mathcal{R}^{n_{\mathbf{features}}}$. There exists some subspace $\mathbf{H}$ that contains each of these vectors that is of dimension at most $n_{\mathbf{classes}}-1$. Directions in the data orthogonal to $\mathbf{H}$ contribute to each class equally and thus do not contribute in distinguishing classes. Thus, computing Euclidean distances in $\mathbf{H}$ will allow us to reduce the data to dimension $n_{\mathbf{classes}}-1$ without sacrificing any information useful for classification. To obtain a lower-dimensional latent space optimal for classification, one can project to a linear subspace $\mathbf{H_{L}}$ that maximizes the variance of the class means after projection. $\mathbf{H_{L}}$ can be obtained through performing PCA on the class means. LDA dimensionality reduction can be useful because it will preserve the inter-class variance in the data, which is useful in discriminating between classes. However, in doing so, the intra-class variance is lost. Thus, understanding relationships between points in the same class can be difficult.
A.3 SHAP Values and Saliency Maps
@@
$\mathbf{SHAP}_{i}=\sum_{S\subseteq F\setminus i}\frac{|S|!(n_{\mathbf{features}}-|S|-1)!}{n_{\mathbf{features}}!}\ E[f(z)|z_{S}]-E[f(z\setminus i)|z_{S\setminus i}]$ (14)
In Eq. 14, $S$ represents a subset of the feature set $F$ that does not include $i$, and $z_{S}$ represents a vector of features including $S$. SHAP is useful because it can assign quantitative, robust feature importance values for each feature regardless of model architecture, which can then be used to either interpret model predictions or perform feature selection. However, SHAP scales poorly when $n_{\mathbf{features}}$ becomes large[15], and the SHAP values can become less valuable for identifying structure-property relationships when using features that do not carry physical meaning. [13]

### v2

@@ B.1 Linear and Kernel Principal Components Analysis
$\ell_{proj}=||\mathbf{X-TP}_{TX}||^{2}/n_{\text{{samples}}}$ (6)
@@
B.2 Linear Discriminant Analysis
@@
$f_{k}(x)=\frac{1}{(2\pi)^{n_{\text{{samples}}}/2}|\Sigma_{k}|^{1/2}}e^{-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})}$ (7)
@@
For dimensionality reduction, LDA will identify a linear projection that maximizes the separation between classes. One can consider the $n_{\text{classes}}$ class means as vectors in $\mathcal{R}^{n_{\text{features}}}$. There exists some subspace $\mathbf{H}$ that contains each of these vectors that is of dimension at most $n_{\text{classes}}-1$. Directions in the data orthogonal to $\mathbf{H}$ contribute to each class equally and thus do not contribute in distinguishing classes. Thus, computing Euclidean distances in $\mathbf{H}$ will allow us to reduce the data to dimension $n_{\text{classes}}-1$ without sacrificing any information useful for classification. To obtain a lower-dimensional latent space optimal for classification, one can project to a linear subspace $\mathbf{H_{L}}$ that maximizes the variance of the class means after projection. $\mathbf{H_{L}}$ can be obtained through performing PCA on the class means. LDA dimensionality reduction can be useful because it will preserve the inter-class variance in the data, which is useful in discriminating between classes. However, in doing so, the intra-class variance is lost, restricting the ability to understanding relationships between points in the same class.
B.3 SHAP Values and Saliency Maps
@@
$\mathbf{SHAP}_{i}=\sum_{S\subseteq F\setminus i}\frac{|S|!(n_{\text{features}}-|S|-1)!}{n_{\text{features}}!}\ E[f(z)|z_{S}]-E[f(z\setminus i)|z_{S\setminus i}]$ (9)
In Eq. (9), $S$ represents a subset of the feature set $F$ that does not include $i$, and $z_{S}$ represents a vector of features including $S$. SHAP is useful because it can assign quantitative, robust feature importance values for each feature regardless of model architecture, which can then be used to either interpret model predictions or perform feature selection. However, SHAP scales poorly when $n_{\text{features}}$ becomes large [15], and the SHAP values can become less valuable for identifying structure-property relationships when using features that do not carry physical meaning. [13]